### v1

@@ 4.3 Finding the Optimal Attributions for Vision Transformer
Figure 5: Attribution-based pruning using a different number of reference samples (per class) to estimate the importance of attention heads (left) or neurons in linear layers (right) of the ViT-B-16. This experiment has been conducted for 20 different random seeds. For the propagation of LRP, LRP-$\epsilon$ has been set as our parameter for all layers (Sec. 3.3), and w.r.t. the attribution of softmax operations (Sec. 0.A.4).
Figure 6: Pruning a ViT model pre-trained on ImageNet (simplified task to detecting three classes), using ten reference samples per class. Results show a better sparsification-performance tradeoff for our optimized LRP composite compared to a heuristic (faithful) LRP composite, Yeom et al. [49] (details for each in Tab. A.2, Sec. 0.A.5, and Sec. 0.A.3.3, respectively), LRP-$\epsilon$, and random pruning.
@@
There are two structures of interest in the Vision Transformers for pruning in literature: attention heads and linear layers [45, 27]. As observed in Sec. 4.1, the ViT model shows a higher degree of over-parameterization, possibly due to the abundance of neurons in the linear layers, ultimately highlighting the possible value for pruning. We now re-investigate the number of necessary reference samples to robustly estimate the relevance of a model component of a ViT-B-16 model. Interestingly, unlike CNNs, according to Fig. 5, there is no deviation in pruning reliability and stability by using different numbers of reference samples. Consequently, our experiments have been applied with exact same settings as in Sec. 4.2.
Our results from Fig. 6 again confirm an improvement of our optimized attribution-based pruning scheme (best LRP composites shown in Tab. A.1) compared to the previous work of [49]. Adopting the generally reliable composite of LRP-$\epsilon$ obtained from CNNs (as in Fig. 4 and Tab. A.2) is again a promising option. Nonetheless, unlike for CNNs, a recently proposed faithful LRP composite [3] designed for the ViT-B-16 model (detailed in Sec. 0.A.5), is not ideal for pruning. This finding, and our experiment in Sec. 0.F.2 show that optimizing an attributor for two different contexts of faithfulness (input or latent space/pruning) (Sec. 3.3) does not necessarily lead to an attributor that attributes faithfully in both input and latent space. The poorer performance of heuristically tuned LRP from [49] (Sec. 0.A.3.3) compared to random when pruning attention heads, underscores both the greater challenge of pruning this structure and the significance of our proposed optimization procedure.
@@ 5 Conclusion
Appendix 0.A Attribution Methods
@@ 0.A.2 Variants of LRP
0.A.3 Tackling Linear and Convolution Layers
@@ 0.A.3.4 LRP-$\gamma$
Evidently, in case of $\gamma=\infty$, LRP-$\gamma$ will be equivalent to LRP-$z^{+}$. The value of $\gamma$ can also be tuned in the optimization process. However, to prevent further computational costs unlike results from [2] in Sec. 0.A.5, we refrain from tuning it.
0.A.4 Tackling Attention and Softmax Non-linearity
@@ 0.A.5 Faithful LRP Composites
Appendix 0.B Spatial Dimension of To-Be-Pruned Components
@@
Appendix 0.C Additionally on Overparameterization
@@
Figure A.1: With settings similar to Fig. 3, few-shot pruning demonstrates overparameterization exceeds, especially in cases of simpler tasks. Additionally, Relevance denoted in the legend is computed by LRP framework with $\epsilon$ rule (see Sec. 0.A.2) applied across all layers.
Appendix 0.D On Optimization: From Bayesian to Grid Search
@@
Appendix 0.E On Number of Reference Samples
@@
Appendix 0.F Optimized Composites
0.F.1 Top Composites for Pruning
@@
0.F.2 Evaluating the Composite
As described in Sec. 3.3, we proposed our pruning framework as a means of evaluating attribution methods in terms of aligning with the actual importance of latent components. However, it is noted that composites highly scored by this criterion, are not necessarily performant in generating explanation heatmaps. Our results in Sec. 4.3 exhibited that LRP composite (Sec. 0.A.5) suggested by [2] is not the most suitable option for pruning. However, qualitative and quantitative experiment (based on [6, 5, 2]) in Fig. A.3 demonstrates that our best composite (see Tab. A.1) is not proper to render explanation heatmap. All in all, LRP explainers tuned to perform well in pruning, do not necessarily induce reliable heatmaps.
@@ 0.F.2 Evaluating the Composite
Appendix 0.G Layer-wise Relevance Flow
@@
Appendix 0.H Change in Behaviour
@@
Figure A.10: Changes in ResNet-18 decision-making for domain restriction of 3 randomly chosen classes (Box Turtle, Alligator Lizard, and Moving Van) via LRP heatmaps (using composite described in Sec. 0.A.5) indicate model failure in high compression rates are caused by high influence of surrounding objects and background despite considering the main object itself. Heatmaps of the model with high pruning rates demonstrate an effect caused by a subsampling shortcut connection becoming apparent.
Figure A.11: Changes in ResNet-50 decision-making for domain restriction of 3 randomly chosen classes (Flamingo, Gordon Setter, and Snowplow) via LRP heatmaps (using composite described in Sec. 0.A.5) indicate model failure in high compression rates are caused by high influence of surrounding objects and unlike ResNet-18 ignores the main object. The shortcut connections of ResNet architecture result in checkerboard pattern exhibited in the heatmaps.
Figure A.12: Changes in VGG-16-BN decision-making for domain restriction of 3 randomly chosen classes (Digital Clock, Patio, Terrace, and Prayer Rug) via LRP heatmaps (using composite described in Sec. 0.A.5) indicate lower pruning capabilities of this architecture compared to ResNet-18 and ResNet-50.

### v2

@@ 4.3 Finding the Optimal Attributions for Vision Transformer
Figure 5: Attribution-based pruning using a different number of reference samples (per class) to estimate the importance of attention heads (left) or neurons in linear layers (right) of the ViT-B-16. This experiment has been conducted for 20 different random seeds. For the propagation of LRP, LRP-$\epsilon$ has been set as our parameter for all layers (Sec. 3.3), and w.r.t. the attribution of softmax operations (Sec. A.4.2).
Figure 6: Pruning a ViT model pre-trained on ImageNet (simplified task to detecting three classes), using ten reference samples per class. Results show a better sparsification-performance tradeoff for our optimized LRP composite compared to a heuristic (faithful) LRP composite, Yeom et al. [48] (details for each in Tab. A.2, Sec. A.5, and Sec. A.3.3, respectively), LRP-$\epsilon$, Integrated Gradient, and random pruning.
@@
Attention heads and linear layers are two structures of interest in the pruning literature for Vision Transformers [44, 26]. As observed in Sec. 4.1, the ViT model shows a higher degree of over-parameterization, possibly due to the abundance of neurons in the linear layers, ultimately highlighting the possible value for pruning. We now re-investigate the number of reference samples required to robustly estimate the relevance of a component in the ViT-B-16 model. Interestingly, unlike CNNs, according to Fig. 5, there is no deviation in pruning reliability and stability by using different numbers of reference samples. Consequently, our experiments have been applied with exact same settings as in Sec. 4.2.
Our results from Fig. 6 again confirm an improvement of our optimized attribution-based pruning scheme (best LRP composites shown in Tab. A.1) compared to the previous work of [48]. Adopting the generally reliable composite of LRP-$\epsilon$ obtained from CNNs (as in Fig. 4 and Tab. A.2) is again a promising option. Nonetheless, unlike for CNNs, a recently proposed faithful LRP composite [2] designed for the ViT-B-16 model (detailed in Sec. A.5), is not ideal for pruning. This finding, and our experiment in Sec. F.2 show that optimizing an attributor for two different contexts of faithfulness (input or latent space/pruning) (Sec. 3.3) does not necessarily lead to an attributor that attributes faithfully in both input and latent space. The poorer performance of heuristically tuned LRP from [48] (Sec. A.3.3) compared to random when pruning attention heads, underscores both the greater challenge of pruning this structure and the significance of our proposed optimization procedure.
@@ 5 Conclusion
Acknowledgements
This work was supported by the Federal Ministry of Education and Research (BMBF) as grant BIFOLD (01IS18025A, 01IS180371I); the German Research Foundation (DFG) as research unit DeSBi (KI-FOR 5363); the European Union’s Horizon Europe research and innovation programme (EU Horizon Europe) as grant TEMA (101093003); the European Union’s Horizon 2020 research and innovation programme (EU Horizon 2020) as grant iToBoS (965221).
@@ Acknowledgements
Appendix A Attribution Methods
@@ A.2 Variants of LRP
A.3 Tackling Linear and Convolution Layers
@@ A.3.4 LRP-$\gamma$
Evidently, in case of $\gamma=\infty$, LRP-$\gamma$ will be equivalent to LRP-$z^{+}$. The value of $\gamma$ can also be tuned in the optimization process. However, to prevent further computational costs unlike results from [2] in Sec. A.5, we refrain from tuning it.
A.4 Tackling Non-linearities
A.4.1 Activation Functions
In LRP framework, a non-linear activation function $\sigma$ (e.g., ReLU), with $\sigma_{i}(z_{i})=a_{i}$ is handled similar to Eq. 6, where $z_{ij}=\delta_{ij}z_{i}$ and $\delta$ represents the Kronecker delta.
A.4.2 Attention and Softmax on Transformers
@@ A.5 Faithful LRP Composites
Appendix B Spatial Dimension of To-Be-Pruned Components
@@
Appendix C Additionally on Overparameterization
@@
Figure A.1: With settings similar to Fig. 3, few-shot pruning demonstrates overparameterization exceeds, especially in cases of simpler tasks. Additionally, Relevance denoted in the legend is computed by LRP framework with $\epsilon$ rule (see Sec. A.2) applied across all layers.
Appendix D On Optimization: From Bayesian to Grid Search
@@
Appendix E On Number of Reference Samples
@@
Appendix F Optimized Composites
F.1 Top Composites for Pruning
@@
F.2 Evaluating the Composite
As described in Sec. 3.3, we proposed our pruning framework as a means of evaluating attribution methods in terms of aligning with the actual importance of latent components. However, it is noted that composites highly scored by this criterion, are not necessarily performant in generating explanation heatmaps. Our results in Sec. 4.3 exhibited that LRP composite (Sec. A.5) suggested by [2] is not the most suitable option for pruning. However, qualitative and quantitative experiment (based on [5, 4, 2]) in Fig. A.3 demonstrates that our best composite (see Tab. A.1) is not proper to render explanation heatmap. All in all, LRP explainers tuned to perform well in pruning, do not necessarily induce reliable heatmaps.
@@ F.2 Evaluating the Composite
Appendix G Layer-wise Relevance Flow
@@
Appendix H Change in Behaviour
@@
Figure A.10: Changes in ResNet-18 decision-making for domain restriction of 3 randomly chosen classes (Box Turtle, Alligator Lizard, and Moving Van) via LRP heatmaps (using composite described in Sec. A.5) indicate model failure in high compression rates are caused by high influence of surrounding objects and background despite considering the main object itself. Heatmaps of the model with high pruning rates demonstrate an effect caused by a subsampling shortcut connection becoming apparent.
Figure A.11: Changes in ResNet-50 decision-making for domain restriction of 3 randomly chosen classes (Flamingo, Gordon Setter, and Snowplow) via LRP heatmaps (using composite described in Sec. A.5) indicate model failure in high compression rates are caused by high influence of surrounding objects and unlike ResNet-18 ignores the main object. The shortcut connections of ResNet architecture result in checkerboard pattern exhibited in the heatmaps.
Figure A.12: Changes in VGG-16-BN decision-making for domain restriction of 3 randomly chosen classes (Digital Clock, Patio, Terrace, and Prayer Rug) via LRP heatmaps (using composite described in Sec. A.5) indicate lower pruning capabilities of this architecture compared to ResNet-18 and ResNet-50.